# Thinking twice inside the box: is Wigner’s friend really quantum?

### v1

@@ II.3 When classically duplicated agents reason about each others’ reasoning
  • (Q): Quantum theory is universally valid.
  • (C): The predictions of different agents must be consistent.
@@
Thought Experiment 3. Imagine Freya is to be put to sleep, and multiplied into $N$ copies. For each copy, now asleep in $N$ different labs, a fair coin is tossed. In each case, if the outcome is Heads, the copy of Freya is duplicated again, or if the outcome is Tails, she is not. Then, each copy of Freya is woken and asked to give her credence that the outcome of her lab’s coin toss was Tails. (We assume that she cannot notice the presence/absence of an identical copy of herself in the lab). This scenario is sketched in Figure 3.
@@
In fact, Freya is offered a bet: she can buy a ticket from a bookie for $(2/3-\varepsilon)$\$, where $\varepsilon>0$ is small, (say, for $66$ cents) that wagers on the coin toss having shown Heads. Meanwhile, superobserving Wigner, outside a certain lab, is offered the same opportunity. It is natural to argue that Freya should buy the ticket, but Wigner should not. That is, the credence that Freya should assign to Tails (which directly determines the maximum price $1-p$ she rationally ought to be prepared to pay) is $1/3$, whilst for Wigner it is $1/2$. Finally, all copies of Freya survive the experiment and are released. Everyone who has bought the ticket now receives $1$\$ if the outcome of their lab’s coin toss was indeed Tails. Freya and Wigner have been initially informed about all the details of the experiment.
@@
Figure 3: Sketch of the setup of Thought Experiment 3; Freya agrees to an experiment in which she will be put to sleep and multiplied into $N$ copies, where $N$ is large (here though, $N=2$). For each copy, asleep in her own lab, a fair coin is tossed. If the outcome is Heads, the copy of Freya in the corresponding lab is duplicated again. If the outcome is Tails, she is not.
@@
Given that half as many copies of Freya will experience waking under the outcome Tails than the outcome Heads, a given copy of Freya may assign a proportionately lower degree of belief that she is contained in the smaller group of copies. There is significant disagreement in philosophical literature concerning whether the original Sleeping Beauty ought to be a “Halfer” [37] or a “Thirder” [38], and depending on how you operationalise the setup [39], one may take either position. Here, we follow the prescription of Elga [40] for the specific question asked in our Thought Experiment: Elga would claim that Freya should assign a uniform probability distribution to self-locating as any copy (see however our discussion in Subsection III.3, where we will reevaluate this option more explicitly, and see Appendix V.4 for an overview on Elga’s principle). Using the law of large numbers, there will almost surely be approximately $N$ copies of Freya who will experience a Heads-awakening, following the duplication process, whilst approximately $N/2$ will experience a Tails-awakening (to first order in $N$). Hence, with high probability, approximately $N/2$ copies of Freya will lose their wager of $(2/3-\epsilon)$\$, whilst approximately $N$ copies of Freya will have profited by $(1/3+\epsilon)$\$. The bet is therefore rational for any copy of Freya to accept for any $\varepsilon>0$ (but not for negative $\varepsilon$) – thus setting her credence for Tails as $1/3$. On the other hand, superobserving Wigner outside a given lab, in his conviction that a fair coin was tossed, should place the probability of Tails as $p=1/2$, which is the threshold price under which he wins on average. Accordingly, we see a divergence, or an apparent inconsistency, in the credences that Freya and Wigner ought to assign to the same event, despite both being ideally rational agents, and having the same objective knowledge about the world.
@@
As in the original Sleeping Beauty problem, the puzzle can be made more dramatic by imagining that Freya will be multiplied into $M$ copies if the outcome of the coin toss is Heads, where $M$ is large (above, we have discussed the case $M=2$). In this case, Freya’s credence in the outcome Tails would accordingly be argued to be $1/(M+1)$, which is vanishingly small, whilst Wigner would still surely assign probability $1/2$.
It is worth immediately noting that this is not related to Ego Absolutism, as the authors of [24] note that “a thought can be absolutely real even if it corresponds to an incorrect statement”. In our example, whether Freya’s probability assignments and beliefs are correct are tangential to whether they exist absolutely. However, the above highlights another kind of inconsistency. Despite both being ideally rational, and despite both having precisely the same knowledge about the world, Freya and Wigner assign different credences to the outcome of the coin toss – or at least they should from the perspective of winning bets, following Elga’s principle. That is, Thirders would argue that Freya’s knowledge that there are many more copies of herself who experience a “Heads-awakening” should increase her credence about self-locating in one of these labs. Whereas Wigner, in his third-person perspective, should not change his beliefs according to the possible existence of multiple copies of Freya. Here we see the first- or third-person perspectives of agents play a fundamental role in their beliefs, despite both agents having the same absolute knowledge.
An analogy can be drawn to the Frauchiger-Renner Gedankenexperiment. Consider the following probabilistic generalisation of Frauchiger and Renner’s assumption (C):
@@
Assumption (CP). Suppose that agent $A$ has established that “I am certain that agent $A^{\prime}$, upon reasoning within the same theory as the one I am using, and having the exact same knowledge of the world as I, is pretty sure that $x=\chi$ at time $t$.” Then agent $A$ can conclude that “I am pretty sure that $x=\chi$ at time $t$.”
We argue that (CP) is violated for Thought Experiment 3. First, we have already argued that Freya ought to set proportionately higher credence that the coin toss landed on Heads, given the context of her being awoken. Wigner, in knowing the full setup of the experiment, and knowing Freya to be a rational gambler, can also reason that Freya will set her credence to be pretty sure of the outcome Heads. Therefore, following (CP), he too should be pretty sure that the outcome was Heads, and should be prepared to bet accordingly. Moreover, “pretty sure” can be made arbitrarily close to “certain”, as in the original wording of the Assumption (C), by taking the number of Heads-duplications $M$ of Freya to be very large. However, as we have argued, Wigner’s credence that the coin toss landed on Heads ought to be $1/2$, therefore the assumption runs us into a contradiction for our scenario.
The proposed violation of (CP) is in a similar vein to the previous violations of AOE and Ego Absolutism that were presented in Thought Experiments 1 and 2; it is not necessarily manifest, but rather in the ambiguity of the wording. The question can be asked differently, such that the observer-dependence is made explicit: “Will I find myself in a world where the coin shows Heads?”. The transitivity of knowledge that is assumed by (C) and (CP) is therefore no longer relevant; although Wigner may reason “I am certain that Freya, upon reasoning within the same theory as the one I am using, and having the same exact knowledge of the world as I, is pretty sure that she will find herself in a world where $c=H$ at time $t$”, this is not to say that Wigner should conclude “I am pretty sure that $c=H$ at time $t$”. It is not even to say that Wigner should believe “I am pretty sure that Freya will find herself in a world where $c=H$ at time $t$”, as this is not demanded by the assumption of Consistency, which refers just to measurable properties.
Moreover, we can resolve the problem (which more closely resembles the conundrum posed by the original Sleeping Beauty problem) of Freya’s credences changing despite an apparent absence of new evidence in a similar fashion. Previously, we imagined that Freya is asked about the outcome of the coin toss, and resolved that Freya would assign credence $1/2$ prior to the duplication experiment, but $1/3$ afterwards. Now we can rephrase the question as one of two, more specific, agent-(in)dependent formulations, to make clear exactly what Freya is asking. Prior to the experiment, Freya may either ask “What credence should I assign to the coin toss landing on Heads?” (to which she should answer $1/2$), or she may ask a more relevant question for the scenario “Under the condition that somebody, later, is physically and psychologically continuous with me, what credence should I assign to the possibility that that person will find themself in a world where the coin has landed on Heads?” (to which she should answer $2/3$). Clearly the latter is overly verbose, so it is unsurprising (but nevertheless misleading) that this is typically abbreviated as in the original problem. Still, the question should be understood to contain an implicit agent-dependence, which is the source of the inconsistency between the credences of varying agents, and over varying times.
@@ II.4 Resource costs of implementation on a classical computer
[…]
Let us examine the specific costs for the operations involved in our thought experiments.
Fission. If quantum phenomena play no significant role in brain processing, then the no-cloning theorem imposes no restrictions on duplication, since it does not apply to orthogonal, classical states. We therefore consider a classical copy operation $(a,0)\mapsto(a,a)$ by the function
$f:(a,b)\mapsto(a,a\oplus b),$
where the initial “ready” state of the bits will be $b=0$. Clearly this requires twice the storage and processing capacity of simulating a single mind, as well as an additional $S$ “copy” gates to transfer the data of the existing mind to a second.
Fusion. To recombine the two observers, we consider this in two steps. First, we need to reverse any changes that had occurred since the duplication (memory erasure), for the observers to be qualitatively identical once again. Then, either we could consider deleting one copy (this would be sufficient for e.g. Parfit’s account of personal identity), or we could construct a more elaborate “merge” operation – for example, taking all of the even bits from one copy, and all of the odd bits from the other. The latter would perhaps be necessary to satisfy those who posit something further to identity than just psychological similarity (and/or physical similarity, for non-computer-generated minds). However, for now we will assume we are just concerned with the abstract information-theoretic content. For the first step, we can look to the reversibility protocol of [24]. They assume that reversibility can be naively achieved by replacing each gate with a reversible one using ancillae bits (i.e. Toffoli gates [47]):
$g:(a,b,c)\mapsto(a,b,c\oplus ab).$
This results in a space overhead that is linear in the number of gates required for brain emulation: $s_{rev}=s+g$ bits (i.e. one additional bit for each gate). This brings the requirements to respective orders of magnitude $s_{rev}=10^{19}$, $g_{rev}=10^{19}$, $t_{rev}=10^{11}$. Now, with two qualitatively identical copies, we could simply clear the data of one copy, using e.g.
$h:(a,b)\mapsto(a,0),$
leaving one copy who believes itself to be and qualitatively is the same observer as the one prior to fission. More elaborately, we could consider an operation in which corresponding bits are compared from each copy, and one of the two deleted if they are identical. After iterating through each bit, this eventually leaves just one complete copy, with an established causal dependency on both prior copies. This is discussed in some more depth in [28]. By either construction though, the resource requirements are of the same order of magnitude as that required for a reversible AI, as estimated in [24].
By contrast, implementing such an AI on a quantum computer, the estimates of [24] must account for the additional logical and physical quantum resources, as well as fault tolerant quantum gates and error correction. Fault tolerant quantum gates are much slower (7 orders of magnitude) than classical gates. This, combined with the overhead due to gate synthesis (3 orders of magnitude), entails that the time for the AI to have thoughts is of order $T_{Q}\sim 10^{10}$ s (or 500–600 years). Given that the initial target was $T=1$ s, the feasibility of the quantum experiment requires major advances in quantum computing – possibilities for which the authors discuss in section 6.6 of [24]. Meanwhile, the classical implementation of such an AI, with its advantage of at least 7 orders of magnitude, will clearly be feasible far sooner.
@@ II.5 Are these fair analogies?
Irrespective of the relative implementability of the quantum and classical experiments, the skeptical reader might be asking, more fundamentally: Why should we be interested in these analogies at all? For most (non-Everettian) quantum theorists, the thought experiments we present here are metaphysically very different from what we may think of as happening during Wigner’s friend scenarios. Why then should we think of them as exhibiting any kind of common features? The skeptic may even accept at this stage that Ego Absolutism (or an analogous assumption) fails to hold in both sets of cases, but maintain that the core feature, respectively picked out by the classical and quantum scenarios, is something fundamentally different.
@@
It is worth noting the disanalogy between our article and the toy model program of Spekkens [22], in that we are not trying to reproduce quantum phenomena classically, but rather features, or alleged consequences, of quantum theory. In this sense, our goal is perhaps more closely related to the research program of Del Santo and Gisin [51, 52, 53], who similarly tackle the question of which features are unique to or characteristic of quantum physics, with their focus on indeterminism in classical physics (including a classical analogue of the measurement problem via objective, ontological indeterminacy).
@@
In the same vein, we contend that the supposedly “quantum” feature captured by Wigner’s friend and the subsequent no-go theorems is not unique to quantum physics. Of course, we do not argue that classical variants of EWF scenarios can reproduce the quantum phenomenology (in particular, by violating an LF inequality): this necessarily involves the violation of a Bell inequality, which is provably incompatible with any classical explanation (in the sense of a local hidden variable model). However, the novelty of the experiment was certainly not the violation of Bell-type inequalities, which is already well established. The novel claim instead is the possibility for violating a combination of (three or four) metaphysical assumptions – and the most interesting of those, we argue, are already untenable for analogous classical experiments. Therefore, to formulate our claim in a more precise way, we have to give a precise definition of the feature that we believe is captured by the new proposed experiments, and which can be classically reproduced. In Section III, we will do so, and define a property termed “Restriction A”, capturing the impossibility in some scenarios of obtaining an absolute probability space for one or more agents from a given physical theory – which we see as the common core of Wigner’s friend and other classical puzzles appearing in physics and philosophy.
@@
For Everettians though, our classical thought experiments can be seen as more directly analogous with their quantum counterparts, as Wigner’s friend already contains some kind of duplication of persons. In particular, Wallace [33] remarks on the philosophical challenges for personal identity in the many-worlds interpretation [54], arguing that branching forces us towards a theory of identity in which multiple persons supervene on a single state. This attitude can be discerned in physics literature too, such as [4]. Deutsch writes that the formalism of quantum theory is inconsistent with there being an “actual” result of a measurement, distinguished from other possible outcomes – presenting a tension with our experience, which the MWI addresses by positing that observables are typically multivalued, possessing all eigenvalues across a multiplicity of worlds. Applying this proposal to Wigner’s friend entails that the friend’s interaction with a quantum system described by a cat state creates two copies of the friend, each of whom observe one of the two outcomes. An Everettian account of Wigner’s friend is therefore metaphysically very similar to our Thought Experiment 1 – only our persons are side-by-side, rather than displaced across worlds. Furthermore, [4] proposes an interference experiment involving agents, to test the predictions of many-worlds against collapse interpretations. The observation of interference effects is said to allow the agent to infer that “there was more than one copy of [themself] (and the atom) in existence” at the time of the measurement, and that “these copies merged to form [their] present self”. At this point, the agents are now said to be once again identical, despite having previously been in two different branches – in precise analogy with our Thought Experiment 2.
@@
This being said, our claim is certainly not ontological in nature – in particular, we do not try to give an account of what actually happens inside the box during a Wigner’s friend experiment. Rather, we claim that the similarities between our thought experiments and certain interpretations of quantum experiments (and, in particular, in their metaphysical consequences) motivate us to ask whether there is a common feature of general physical theories that characterises these puzzles. This leads us to ask about the common, interpretation-independent, structural and mathematical elements of such scenarios, motivating the following section.

### v2

@@ II.3 When classically duplicated agents reason about each others’ reasoning
  • (Q): Quantum theory is universally valid;
  • (C): The predictions of different agents must be consistent;
@@
Thought Experiment 2. Imagine Freya and Wigner are to be put to sleep, and multiplied into $N$ copies. Each couple is distributed to one of $N$ identical laboratories. In each laboratory, a fair coin is tossed, and if the outcome is Heads, the copy of Freya (but not the copy of Wigner) is duplicated again. Then, all participants are woken and asked to give their credence that the outcome of their lab’s coin toss was Tails. (We assume that they cannot notice the presence/absence of an identical copy of Freya in the lab). This scenario is sketched in Figure 3.
@@
In fact, all participants are offered a bet: they can buy a ticket from a bookie for $(2/3-\varepsilon)$\$, where $\varepsilon>0$ is small, (say, for $66$ cents) that wagers on the coin toss having shown Heads. It is natural to argue (see below) that the credence that Freya should assign to Tails (which directly determines the maximum price $1-p$ she rationally ought to be prepared to pay) is $1/3$, whilst for Wigner it is $1/2$: Freya should buy the ticket, but Wigner should not.
Finally, all copies survive the experiment and are released. Everyone who has bought the ticket now receives $1$\$ if the outcome of their lab’s coin toss was indeed Tails. Freya and Wigner have been initially informed about all the details of the experiment.
@@
Figure 3: Sketch of the setup of Thought Experiment II.3; Freya and Wigner agree to an experiment in which they will be put to sleep and multiplied into $N$ copies, where $N$ is large (here though, $N=2$). For each lab, a fair coin is tossed. If the outcome is Heads, the copy of Freya (but not of Wigner) in the corresponding lab is duplicated again. If the outcome is Tails, she is not.
@@
Given that half as many copies of Freya will experience waking under the outcome Tails than the outcome Heads, a given copy of Freya may assign a proportionately lower degree of belief that she is contained in the smaller group of copies. There is significant disagreement in philosophical literature concerning whether the original Sleeping Beauty ought to be a “Thirder” [38] or a “Halfer” [39], and depending on how you operationalise the setup [40], one may take either position. Here, we follow the prescription of Elga [41] for the specific question asked in our Thought Experiment: Elga would claim that Freya should assign a uniform probability distribution to self-locating as any copy (see Appendix V.4 for an overview on Elga’s Principle of Indifference [38, 41], and Subsection III.3, in which we reevaluate this option more explicitly). Using the law of large numbers, there will almost surely be approximately $N$ copies of Freya who will experience a Heads-awakening, following the duplication process, whilst approximately $N/2$ will experience a Tails-awakening (to first order in $N$). Hence, with high probability, approximately $N/2$ copies of Freya will lose their wager of $(2/3-\epsilon)$\$, whilst approximately $N$ copies of Freya will have profited by $(1/3+\epsilon)$\$. The bet is therefore rational for any copy of Freya to accept for any $\varepsilon>0$ (but not for negative $\varepsilon$) – thus setting her credence for Tails as $1/3$. Similarly, about $N/2$ copies of Wigner will lose their wager, while approximately $N/2$ copies will have profited. Accordingly, Elga’s principle implies the following diverging credences for Freya and Wigner respectively:
$P_{F}(T)\approx\frac{N/2}{N+N/2}=\frac{1}{3},\qquad P_{W}(T)\approx\frac{N/2}{N/2+N/2}=\frac{1}{2},$ (2)
where $P_{F}(T)$ is given by the proportion of “Tails-Freyas” (Freyas who will see the outcome Tails) to the total number of Freyas, and likewise for $P_{W}(T)$.
@@
As in the original Sleeping Beauty problem, the puzzle can be made more dramatic by imagining that Freya will be multiplied into $M$ copies if the outcome of the coin toss is Heads, where $M$ is large (above, we have discussed the case $M=2$). In this case, Freya’s credence in the outcome Tails would accordingly be argued to be $1/(M+1)$, which is vanishingly small, whilst Wigner would still surely assign probability $1/2$:
$P_{F}(T)\approx\frac{N/2}{NM/2+N/2}=\frac{1}{M+1}\stackrel{{\scriptstyle M\to\infty}}{{\longrightarrow}}0,\qquad P_{W}(T)\approx\frac{1}{2}.$ (3)
As indicated by the $\approx$” signs, the analysis above is not yet fully rigorous: strictly speaking, Elga’s principle applies only to situations in which there is a fixed, predetermined, known number of copies of an agent. Here, however, the number of copies depends on the number of Heads results of the $N$ coin tosses. With probability approaching unity for large $N$, this number is about $N/2$ (up to plus or minus $\varepsilon N$ for any small $\varepsilon>0$). Hence, regardless of how we apply Elga’s principle to situations with a probabilistically varying number of copies, the above equations should become exact in the limit of $N\to\infty$.
For what follows, however, we will be interested in finite $N$ (in fact, $N=1$). To this end, we need to implement one further assumption:
Assumption (NUI) (no update on irrelevant information). Suppose that Freya knows that she is one of several identical copies in an experiment, and that all copies will undergo the exact same procedure, independently, until the experiment is completed. Then, her predictions for her future observations will not change if she is told which one of the copies she is.
That beliefs should not change in the absence of new evidence is a very intuitive assumption, and it is related to van Fraassen’s reflection principle [42], as remarked by Groisman [40]. Consider a modification of Thought Experiment II.3, where shortly before the coin tosses, each one of the $N$ Freyas is told the number $j\in\{1,\ldots,N\}$ of the laboratory that they are in. By Assumption (NUI), if $N$ is very large, Freya’s probability assignment $P_{F}(T)\approx\frac{1}{3}$ should not change by this. Indeed, consider the copy of Freya in lab $j$ (call her Freyaj). From her perspective, what follows is an experiment with a single lab in which, depending on the outcome of a single coin toss, she will be duplicated or not. That is, Freyaj is subject to Thought Experiment II.3 for the special case $N=1$, and Assumption (NUI) leads her to assign the probability $P_{F_{j}}(T)=\frac{1}{3}$.
In summary, Elga’s principle in conjunction with (NUI) implies the following: if a random process selects one of $n$ possible outcomes with probabilities $\{p_{i}\}_{i=1}^{n}$, and for outcome $i$ she will be copied $m_{i}$ times, then she should assign the probability to observe outcome $i$ proportional to $m_{i}p_{i}$. It is reassuring (and not at all obvious) that this conclusion is consistent with the above analysis.
This can be shown as follows. In Thought Experiment II.3, for an outcome of $k$ Heads from $N$ labs, there will be $Mk$ copies of Freya in Heads labs, and $N-k$ copies in Tails labs — therefore a total number of $N+(M-1)k$ copies of Freya, who will wake up and see the outcome $k$. Since the Binomial probability of this number of Heads is (${N\choose k}/2^{N}$, the above prescription entails that Freya’s subjective probability of the outcome $k$ is
$P_{F}(k)=c\,\frac{N+(M-1)k}{2^{N}}\left(\begin{array}[]{c}N\\ k\end{array}\right),$ (4)
where $c>0$ is the proportionality constant. Then, Elga’s principle dictates that Freya has probability $Mk/(N+(M-1)k)$ of observing Heads. This is because there are $Mk$ Heads-Freyas, and the total number of Freyas is $N+(M-1)k$. Therefore, Freya’s total subjective probability of observing Heads is as follows:
$P_{F}(H)=\sum_{k=0}^{N}P_{F}(k)P_{F}(H|k)=c\sum_{k=0}^{N}\frac{N+(M-1)k}{2^{N}}\left(\begin{array}[]{c}N\\ k\end{array}\right)\frac{Mk}{N+(M-1)k}=c\frac{MN}{2}.$ (5)
Similarly, we obtain
$P_{F}(T)=\sum_{k=0}^{N}P_{F}(k)P_{F}(T|k)=c\sum_{k=0}^{N}\frac{N+(M-1)k}{2^{N}}\left(\begin{array}[]{c}N\\ k\end{array}\right)\frac{N-k}{N+(M-1)k}=c\frac{N}{2},$ (6)
and the constant $c$ is fixed by the requirement $P_{F}(H)+P_{F}(T)=1$, resulting in $P_{F}(H)=M/(M+1)$ and $P_{F}(T)=1/(M+1)$, independently of $N$. That is, equation (3) actually holds exactly. Indeed, this is the only possible result consistent with (NUI) and $P_{F_{j}}=\frac{1}{3}$ claimed above.
But now, consider the special case of $N=1$. In this case, Freya observes a Heads outcome if and only if Wigner observes a Heads outcome. Nonetheless, assuming Elga’s principle and (NUI), Freya and Wigner assign very different probabilities to this single outcome, as we have just shown. An analogy can be drawn to the Frauchiger-Renner Gedankenexperiment. Consider the following probabilistic generalisation of Frauchiger and Renner’s assumption (C):
@@
Assumption (CP) (probabilistic consistency). Suppose that agent $A$ has established that “I am certain that agent $A^{\prime}$, upon reasoning within the same theory as the one I am using, and having the exact same knowledge of the world as I, is pretty sure that $x=\chi$ at time $t$.” Then agent $A$ can conclude that “I am pretty sure that $x=\chi$ at time $t$.”
If $M$ is large, and if we assume Elga’s principle and Assumption (NUI), then Wigner will assign probability $P_{W}(T)=\frac{1}{2}$ to the Tails outcome, whereas Freya will assign $P_{F}(T)=\frac{1}{M+1}$. That is, Freya and Wigner have exactly the same information about the experiment, and they use the exact same theory for prediction. However, whilst Freya is pretty sure that she will observe outcome Heads, Wigner is not. This violates Assumption (CP). While the calculations in equations (5) and (6) show that Elga’s principle and (NUI) are not in obvious contradiction to each other (at least in Thought Experiment II.3), we have proven incompatibility of both together with Assumption (CP):
Theorem 1.
For Thought Experiment II.3, the following three assumptions cannot jointly hold:
  • Elga’s Principle of Indifference;
  • Assumption (NUI) (no update on irrelevant information);
  • Assumption (CP) (probabilistic consistency).
This “classical no-go theorem” will be analysed more thoroughly in Section III.1, in particular following Claim 5.
@@ II.4 Resource costs of implementation on a classical computer
Estimates for the storage costs involved in classically simulating the human brain are typically given anywhere in the range of $10^{9}$ to $10^{28}$ bits, with [25] using $S=10^{15}$ bits as an appropriate space approximation. The required rate is taken to be of the order $F=10^{15}$ FLOP/s (floating point operations per second), or $\gamma=10^{19}$ s-1 (gates per unit time). Positing a high degree of parallelisation, this can be taken as a depth per unit time of $\delta=10^{11}$ s-1. For the AI to have thoughts at a similar rate to a human (of the order $T=1$), the model is taken to be of depth $t=\delta T=10^{11}$.
By contrast, implementing such an AI on a quantum computer, the estimates of [25] must account for the additional logical and physical quantum resources, as well as fault tolerant quantum gates and error correction. Fault tolerant quantum gates are much slower (7 orders of magnitude) than classical gates. This, combined with the overhead due to gate synthesis (3 orders of magnitude), entails that the time for the AI to have thoughts is of order $T_{Q}\sim 10^{10}$ s (or 500–600 years). Given that the initial target was $T=1$ s, the feasibility of the quantum experiment requires major advances in quantum computing – possibilities for which the authors discuss in section 6.6 of [25]. Meanwhile, the classical implementation of such an AI, with its advantage of at least 7 orders of magnitude, will clearly be feasible far sooner. And so will be the implementation of our classical duplication (or multiplication) thought experiments: processes such as fission, fusion (as in Thought Experiment II.2) or the parallel simulation of $M\approx 10$ copies should slow down the classical AI implementation by a constant factor of not much more than 2 orders of magnitude, which makes it still much faster than the quantum implementation of a single AI.
@@ II.5 Are these fair analogies?
Irrespective of the relative implementability of the quantum and classical experiments, the skeptical reader might be asking, more fundamentally: Why should we be interested in these analogies at all? For most (non-Everettian) quantum theorists, the thought experiments we present here are metaphysically very different from what we may think of as happening during Wigner’s Friend scenarios. Why then should we think of them as exhibiting any kind of common features? The skeptic may even accept at this stage that some AOE-like assumption fails to hold in both sets of cases, but maintain that the core feature, respectively picked out by the classical and quantum scenarios, is something fundamentally different.
@@
It is worth noting the disanalogy between our article and the toy model program of Spekkens [23], in that we are not trying to reproduce quantum phenomena classically, but rather features, or alleged consequences, of quantum theory. In this sense, our goal is perhaps more closely related to the research program of e.g. Del Santo and Gisin [52, 53, 54], who similarly tackle the question of which features are unique to or characteristic of quantum physics, with their focus on indeterminism in classical physics (including a classical analogue of the measurement problem via objective, ontological indeterminacy).
@@
In the same vein, we contend that the supposedly quantum feature captured by Wigner’s Friend and the subsequent no-go theorems is not unique to quantum physics. Of course, we do not argue that classical variants of EWF scenarios can reproduce the quantum phenomenology (in particular, by violating an LF inequality): this necessarily involves the violation of a Bell inequality, which is provably incompatible with any classical explanation (in the sense of a local hidden variable model). However, the novelty of the experiment was certainly not the violation of Bell-type inequalities, which is already well established. The novel claim instead is the possibility for violating a combination of metaphysical assumptions – and the most interesting of those, we argue, are already untenable for analogous classical experiments. Of course, there are many formulations of EWF arguments, constructed from a variety of metaphysical assumptions - some of which may be better suited to identifying some fundamentally nonclassical feature (e.g. recently, the authors of [19] have proposed a GHZ–FR paradox, based on consistency amongst only communicated measurements). Nevertheless, our analysis has shown that there exist undeniable similarities between some EWF scenarios and classical thought experimentation – and this commonality should not be ignored. To formulate this claim in a more precise way, in Section III we give a definition of the feature that we believe is captured by EWF experiments, and can be classically reproduced. This property, termed “Restriction 1”, captures the impossibility in some scenarios of obtaining a probability distribution for one or more agents from a given physical theory – which we see as the common core of Wigner’s Friend and other classical puzzles appearing in physics and philosophy.
@@
For Everettians though, our classical thought experiments can be seen as more directly analogous with their quantum counterparts, as Wigner’s Friend already contains some kind of duplication of persons. In particular, Wallace [34] remarks on the philosophical challenges for personal identity in the many-worlds interpretation [55], arguing that branching forces us towards a theory of identity in which multiple persons supervene on a single state. This attitude can be discerned in physics literature too, such as [4]. Deutsch writes that the formalism of quantum theory is inconsistent with there being an “actual” result of a measurement, distinguished from other possible outcomes – presenting a tension with our experience, which the MWI addresses by positing that observables are typically multivalued, possessing all eigenvalues across a multiplicity of worlds. Applying this proposal to Wigner’s Friend entails that the Friend’s interaction with a quantum system described by a cat state creates two copies of the Friend, each of whom observe one of the two outcomes. An Everettian account of Wigner’s Friend is therefore metaphysically very similar to our Thought Experiment II.2 – only our persons are side-by-side, rather than displaced across worlds. Furthermore, [4] proposes an interference experiment involving agents, to test the predictions of many-worlds against collapse interpretations. The observation of interference effects is said to allow the agent to infer that “there was more than one copy of [themself] (and the atom) in existence” at the time of the measurement, and that “these copies merged to form [their] present self”. At this point, the agents are now said to be once again identical, despite having previously been in two different branches – in precise analogy with our Thought Experiment II.2.
@@
This being said, our claim is not ontological in nature – in particular, we do not try to give an account of what actually happens inside the box during a Wigner’s Friend experiment. Rather, we claim that the similarities between our thought experiments and certain interpretations of quantum experiments (and, in particular, in their metaphysical consequences) prompt us to ask whether there is a common feature of general physical theories that characterises these puzzles. This leads us to ask about the common, interpretation-independent, structural and mathematical elements of such scenarios, motivating the following section.